OpenStudy (bigbosssaint21):

What are the steps you would use to graph f(x)= -16x^2 + 60x + 16 X intercepts are (-1/4,0) and (4,0)

OpenStudy (anonymous):

Simplify 100x2-60x - 16 Pulling out like terms : 1.1 Pull out like factors : 100x2 - 60x - 16 = 4 • (25x2 - 15x - 4) Trying to factor by splitting the middle term 1.2 Factoring 25x2 - 15x - 4 The first term is, 25x2 its coefficient is 25 . The middle term is, -15x its coefficient is -15 . The last term, "the constant", is -4 Step-1 : Multiply the coefficient of the first term by the constant 25 • -4 = -100 Step-2 : Find two factors of -100 whose sum equals the coefficient of the middle term, which is -15 . -100 + 1 = -99 -50 + 2 = -48 -25 + 4 = -21 -20 + 5 = -15 That's it

OpenStudy (anonymous):

maybe?

OpenStudy (bigbosssaint21):

how do you apply that to my problem?